OpenStudy (rissa):

I need help with the average rate of change f(t)=9/t ,t=a, t=a+h

OpenStudy (photon336):

\[f(t) = \frac{ 9 }{ t}\] \[(t = a) | t = (a+h)\]

OpenStudy (photon336):

what approach comes to mind?

OpenStudy (rissa):

f(b)-f(a)/b-a

OpenStudy (photon336):

do you know how to take the power rule of differentiation? this is average rate so i'm thinking slope here

OpenStudy (rissa):

no im not sure

OpenStudy (photon336):

another way is this, I guess we could use the definition of a limit here. \[\lim_{h \rightarrow 0} (\frac{ f(a)-f(a+h) }{ a-(a+h) })\]

OpenStudy (rissa):

yeah that is the way im talking about

OpenStudy (rissa):

im not sure how to substitute this ?

OpenStudy (photon336):

\[\frac{ \frac{ 9 }{ a } - \frac{ 9 }{ a+h }}{ h } \]

OpenStudy (rissa):

is it like \[f(9/a+h)-f(9/a)/(a+h)-9/a\]

OpenStudy (rissa):

i thought b went before a ?

OpenStudy (photon336):

you are correct i'm wrong

OpenStudy (photon336):

it would be in the form f(a+h)-f(a)

OpenStudy (photon336):

\[\frac{ f(\frac{ 9 }{ a+h })-f(\frac{ 9 }{ a }) }{ h }\]

OpenStudy (rissa):

is this difference quotient?

OpenStudy (photon336):

@Nnesha clarification?

OpenStudy (photon336):

would the x values at the bottom just be (a+h)-(a) @Nnesha

OpenStudy (rissa):

what happens to 9/t?

OpenStudy (photon336):

my assumption is t is substituted where a is t = a

OpenStudy (rissa):

yeah

Nnesha (nnesha):

what x value ?? btw riss is it for calculus class ?

OpenStudy (rissa):

this is for precalc

OpenStudy (rissa):

question 8

OpenStudy (stephy321):

I need help with this. I know that D and E are correct because there is a infinite discontinuity at 1 and also a vertical asymptote. I'm thinking that C is also not the right answer. Can someone help me? Please and thank you!

Nnesha (nnesha):

ahh i see what you were talking about and in this case x is t so a+h but.. the denominator would not be a+h - a we have to find the common denominator

Nnesha (nnesha):

rissa i can't open that file have you started the limit types q in class yet or not ?

OpenStudy (stephy321):

ooop i put my question in your thing. soooo sorry

OpenStudy (rissa):

ive been taking a summer course but i caught bronchitis and then i couldnt be in the class. so they moved my final however the final is suppose to be for fall and spring students which im not. so there are alot of things i dont know on this exam.

OpenStudy (rissa):

Determine the average rate of change of the function f(t) = 9 t between the values t = a and t = a + h. Simplify your answer.

OpenStudy (rissa):

9/t

Nnesha (nnesha):

alright so i don't think we need to take the limit looks like simple algebra question \[\large\rm \frac{ y_2 -y_1}{ x_2 -x_1 }=average ~rate ~of~change\] in this question x=t so substitute t values into the function to find Y_1 and y_2

Nnesha (nnesha):

as mentioned above n Photon336 comment replace `t` with a and `a+h` \[ \frac{ f(\frac{9}{a+h}) -f(\frac{9}{a} )}{ t_2 - t_1 }\] t_2 = a+h and t=a \[ \frac{ f(\frac{9}{a+h}) -f(\frac{9}{a} )}{ \color{Red}{a+h} - \color{blue}{a} }\]

Nnesha (nnesha):

actually i just figured out that is not the correct way to right this let me correct that it should be \[\frac{ f(\color{Red}{a+h})-f(\color{blue}{a}) }{ \color{Red}{a+h}-\color{blue}{a} }\] at the denominator f(a+h) means t=a+h if a+h is input what would be the output ?

Nnesha (nnesha):

\[f(t)=\frac{9}{t}\] replace that `t` with a what would you get ?? and then replace that `t` with `a+h` what would you get ?

OpenStudy (rissa):

9/a

OpenStudy (rissa):

\[(9/a+h)-(9/a) /a+h-9/a\]

Nnesha (nnesha):

input would be at the denominator so t values would be at the denominator a+h-a like this

Nnesha (nnesha):

\[f(a+h)\] means if input is a+h what would be the output simply substitute t with `a+h` into this function \[\rm f(t)=\frac{9}{\color{Red}{t}}\] \[\rm f(\color{ReD}{a+h})=\frac{9}{\color{Red}{a+h}}\]

Nnesha (nnesha):

now replace f(a+h) with 9/a+h

OpenStudy (rissa):

im trying to send you a photo

OpenStudy (rissa):

of where im at

Nnesha (nnesha):

f(a) when input is a the output would be what? to find out replace t with a into the function 9/t \[\rm f(t) =\frac{9}{t} \] \[\rm f(\color{Red}{a}) =\frac{9}{\color{Red}{a}} \] plugin \[\frac{ y_2 -y_1 }{ x_2 -x_1 } = \frac{ \frac{9}{a+h}-\frac{9}{a} }{ a+h-a }\]

Nnesha (nnesha):

how did you get 9/a at the denominator??

OpenStudy (rissa):

i thought the formula was f(b)-f(a)/b-a

OpenStudy (rissa):

i got 9/a because i thought you replace t with a

Nnesha (nnesha):

yeah replace t with a into the original function f(t)=9/t

Nnesha (nnesha):

f(b)-f(a) over b-a is same as y_2 - y_1 over x_2 -x_a b and a are the input value

OpenStudy (rissa):

im so confused :(

OpenStudy (rissa):

is there anyway that you can send me a photo of the set up and how you answered it?

Nnesha (nnesha):

i already posted the half of the solution

OpenStudy (rissa):

im just confused about the bottom

Nnesha (nnesha):

\[\rm \frac{ \frac{9}{a+h}-\frac{9}{a} }{ a+h-a }\] if t=a then f(a) = 9/a if t=a+h then f(a+h)=9/a+h \[\rm \frac{ \color{Red}{\frac{9}{a+h}-\frac{9}{a} }}{ a+h-a }\] simplify the numerator first find the common denominator

OpenStudy (rissa):

omg I GOT it!! im so sorry it took me so long to figure this out.